# Self-Supervised Enhancement of Forward-Looking Sonar Images: Bridging Cross-Modal Degradation Gaps through Feature Space Transformation and Multi-Frame Fusion

### v2

@@ Abstract
Enhancing forward-looking sonar images is critical for accurate underwater target detection. Current deep learning methods mainly rely on supervised training with simulated data, but the difficulty in obtaining high-quality real-world paired data limits their practical use and generalization. Although self-supervised approaches from remote sensing partially alleviate data shortages, they neglect the cross-modal degradation gap between sonar and remote sensing images. Directly transferring pretrained weights often leads to overly smooth sonar images, detail loss, and insufficient brightness. To address this, we propose a feature-space transformation that maps sonar images from the pixel domain to a robust feature domain, effectively bridging the degradation gap. Additionally, our self-supervised multi-frame fusion strategy leverages complementary inter-frame information to naturally remove speckle noise and enhance target-region brightness. Experiments on three self-collected real-world forward-looking sonar datasets show that our method significantly outperforms existing approaches, effectively suppressing noise, preserving detailed edges, and substantially improving brightness, demonstrating strong potential for underwater target detection applications.
@@
Forward-Looking Sonar, Self-Supervised Learning, Multi-Frame Fusion, Cross-Modal Degradation Gap.
@@ I Introduction
Forward-looking sonar (FLS), with its exceptional underwater imaging capability, has become an indispensable tool in underwater target detection, environmental perception, and autonomous navigation tasks[1, 2, 3, 4, 5]. However, in complex underwater environments, sound waves are affected by factors such as water boundary reflection, multipath interference, and sound absorption by target materials, leading to significant degradation of sonar images. Common issues include speckle noise, uneven brightness, and brightness artifacts[6, 7, 8]. These challenges hinder tasks like target detection and 3D reconstruction, limiting the practical application and development of FLS technology in real-world underwater scenarios[9, 10, 11]. Efficiently enhancing forward-looking sonar images in such environments is a key issue in underwater robotics and marine exploration[12, 13, 14].
@@
Figure 1: (a) Bridging the Cross-Modal Gap. The degradation gap between multispectral remote sensing (RS) and forward-looking sonar (FLS) images is effectively reduced by mapping both modalities into a unified feature space using wavelet scattering transform (WST). (b) Enhanced Results with Our Method. Compared to the overly smoothed baseline, our method preserves clear target contours and distinct foreground-background boundaries.
Sonar image enhancement methods can be divided into traditional and deep learning-based approaches. Traditional methods[15, 16, 17] improve image quality through filtering, contrast adjustment, and gray-level transformations but struggle to enhance severely degraded images due to limited information. Deep learning methods, especially supervised learning[18, 19, 20, 21], have made significant progress by learning the mapping from degraded to high-quality images using a large number of high-quality paired training samples. However, obtaining such paired sonar image data in underwater environments is challenging. As a result, existing research[22, 23] relies heavily on simulated data for training, leading to poor generalization to real-world data. Self-supervised learning, which does not require reference images, has emerged as an effective solution for enhancing low-quality sonar images in real scenarios. However, existing methods[24, 25] have not validated their enhancement effects on more complex targets and do not fully address the cross-modal degradation gap between multispectral remote sensing[26, 27, 28] and forward-looking sonar images, as shown in the upper part of Figure 1(a). While both are affected by speckle noise, their imaging mechanisms and noise characteristics differ significantly. Multispectral remote sensing images[29, 30] are approximated by pinhole imaging, suffer from low contrast and blurring due to uneven illumination and atmospheric attenuation, while forward-looking sonar images are distance-based[1, 2], with multipath effects from water scattering resulting in weak contrast and uneven brightness. Thus, directly transferring models trained on remote sensing images does not work effectively for sonar images, leading to issues such as overly smooth enhancement results, missing details, and insufficient brightness, as shown in the upper part of Figure 1(b).
To address these issues, we propose a self-supervised enhancement method for forward-looking sonar images, based on feature space transformation and multi-frame fusion. First, we use traditional feature extraction techniques (such as HOG, Canny, Haar, and Wavelet features)[31, 32, 33, 15, 34] to map the raw sonar image from a low-quality pixel space to a more robust feature space, reducing the cross-modal degradation gap between remote sensing and sonar images, as shown in the lower part of Figure 1(a). Second, a self-supervised multi-frame fusion framework is adopted to fully utilize complementary information between multiple frames for natural denoising and target enhancement. This eliminates the need for high-quality paired training data, improving the model’s generalization and practical applicability in underwater environments. Compared to previous self-supervised methods, our approach effectively removes noise, preserves target edge details, and significantly enhances brightness in target regions, as shown in the lower part of Figure 1(b).
The main contributions of this paper are summarized as follows:
  • Proposes an effective transfer learning approach for sonar image processing, mapping sonar images to a robust feature space to reduce the cross-modal degradation gap with other domains (e.g., remote sensing images).
  • Designs a self-supervised enhancement framework for forward-looking sonar images based on multi-frame fusion, achieving significant enhancement effects without requiring reference images.
  • Constructs a multi-view forward-looking sonar image dataset with three types of targets made from different materials, exhibiting diverse degradation characteristics, and providing accurate pose parameters for each image. Experiments on this dataset show that our method outperforms existing mainstream methods in terms of visual quality and quantitative metrics.
Figure 2: Overview of the proposed method. The main idea is that the model integrates a feature space transformation module and a multi-frame image fusion network to bridge the cross-modal degradation gap and enhance the quality of forward-looking sonar images through self-supervised learning.The original images, heavily degraded by speckle noise and low target-background contrast, are significantly improved after enhancement, exhibiting greater brightness, sharper target contours, and well-preserved edge details.
@@ II-A Quality Degradation of Forward-Looking Sonar Images and Traditional Enhancement Methods
In addition, these methods incorporate various handcrafted feature extraction techniques, such as HOG, Canny edge detection, Haar-like features, and wavelet scattering transform[31, 32, 33, 15, 34]. By extracting key structural information, these techniques transform images from the pixel domain to a more robust feature space, improving image quality under noisy conditions. We argue that such feature-space transformations effectively reduce cross-modal degradation gaps and provide a practical framework for transferring enhancement models trained in other domains to sonar image processing tasks.
@@ II-B Deep-Learning-Based Sonar Image Enhancement Methods
To reduce reliance on paired data, methods combining traditional transform-domain techniques and deep learning have emerged. For instance, FDBANet [41] introduces a frequency-domain denoising attention module using FFT to convert sonar images to the frequency domain, filtering noise and enhancing high-frequency features. While this method works well for low-resolution images, its enhancement is limited due to the lack of a data-driven learning mechanism. Self-supervised learning[42, 43, 44, 45], however, does not require reference images and directly utilizes real low-resolution sonar images for training. The SEGSID framework[24] enhances local receptive fields and extracts global semantic information, overcoming traditional blind spot network (BSN) limitations in handling speckle noise with strong spatial correlation. Yet, it uses only single-image information, limiting its ability to enhance image resolution. In contrast, multi-frame fusion methods leverage complementary information across frames, naturally removing noise and improving resolution[25]. This method boosts the brightness of small target areas by fine-tuning a pre-trained remote sensing model but does not fully address the cross-modal degradation gap between multispectral remote sensing and sonar images[34]. In regions with complex structures or severe degradation, the enhanced image can become overly smooth, losing edge details and affecting subsequent target detection accuracy.
@@ III-A Framework Overview
As illustrated in Figure 2, the self-supervised forward-looking sonar image enhancement method proposed in this study consists of two main components: the feature space transformation module and the multi-frame image fusion enhancement network.
Both multispectral remote sensing and forward-looking sonar images suffer from speckle noise and low resolution, but a cross-modal degradation gap exists between them[46, 47]. Directly transferring pre-trained models from multispectral data to sonar images often leads to overly smooth, detail-lacking, and low-brightness results. To address this, we designed the feature space transformation module. By applying handcrafted filtering features, such as wavelet scattering transforms, we convert the original sonar images from a low-quality pixel domain into a more robust feature space. This transformation effectively reduces the degradation gap between remote sensing and sonar images, enabling better adaptation of pre-trained models to real sonar data.
The extracted features from both the feature and pixel spaces are concatenated and passed through convolution operations before being fed into the multi-frame image fusion enhancement network. Using self-supervised learning, complementary information from consecutive frames is integrated to enhance image details and brightness. As the sonar data includes accurate pose information, no additional image registration is needed before multi-frame fusion. This framework minimizes dependence on paired supervised data, achieving significantly improved enhancement results on three self-collected real-world forward-looking sonar dataset compared to existing methods.
III-B Feature Space Transformation Module
To bridge the cross-modal degradation gap between multispectral remote sensing and forward-looking sonar images, we designed the feature space transformation module. This module maps the original sonar images from pixel space to a more robust feature space using handcrafted filtering features, which enhances noise suppression and captures structural information, thus narrowing the gap and providing consistent input for the subsequent self-supervised multi-frame fusion enhancement.
Figure 3: Visualization of handcrafted features. WST is selected for its balanced performance in denoising and detail retention (see Ablation Study).
The module applies a series of predefined filters to the original single-channel sonar images $I$, extracting robust feature space information. Figure 3 shows the visual results of these features:
$G_{i}(I)=\Phi_{i}(I),i\in\{\text{HOG},\text{Canny},\text{GRE},\text{Haar},\text{WST}\}$ (1)
where ${\Phi_{i}}$ denotes their respective filtering transformation functions, and ${G_{i}(I)}$ denotes the extracted filtering features.Next, we concatenate the original image $I$ with the extracted features ${G_{i}(I)}$, creating an enhanced input representation:
$I_{aug}=\text{concat}(I,\{G_{i}(I)\})$ (2)
To integrate these multi-channel features, we use $1\times 1$ convolution operations ${\rm{\Psi}}(\cdot)$ to compress and unify the channels, producing the final fused output:
$I_{fused}=\Psi(I_{aug})$ (3)
As shown in Figure 1, the original forward-looking sonar image (red) and multispectral remote sensing image (green) exhibit a cross-modal degradation gap. The feature space transformation module serves as a bridge, converting the data from the heterogeneous pixel space to a unified feature space (yellow), significantly reducing the gap and enabling better adaptation of pre-trained models from remote sensing data to our self-supervised multi-frame fusion network.
@@ III-C Multi-Frame Image Fusion Network
Figure 4: Illustration of the multi-frame fusion network. The proposed network recursively fuses aligned forward-looking sonar frames in a pairwise manner—each step combines the current state with a new frame—until all K frames are aggregated. A median reference image guides self-supervised enhancement through downsampling and gradient consistency losses, enabling effective speckle noise suppression and high-resolution output (from 376×376 to 800×800).
Forward-looking sonar systems typically capture images at high frame rates, which allows for the effective use of complementary information between consecutive frames to naturally eliminate speckle noise. As shown in Figure 4, our designed multi-frame image fusion network does not require an additional registration module because the collected sonar images come with accurate pose information.
The overall architecture comprises three modules: encoding, fusion, and decoding. Let the input low-resolution image sequence be denoted by
$J=\{J(1),J(2),\ldots,J(K)\}$ (4)
where ${J^{(i)}}$ represents the $i^{\text{th}}$ frame. The encoding module ${E(\cdot)}$ extracts features from each frame, mapping the sequence to a high-dimensional feature space; the fusion module ${F(\cdot)}$ integrates these features into a global encoding ${Z=F(E(J))}$ and the decoding module ${D(\cdot)}$ upsamples and reconstructs the global encoding ${Z}$ to generate the final super-resolution image ${Y}$, namely
$Y=D(F(E(J)))$ (5)
Since it is challenging to acquire high-quality sonar images in real underwater scenarios, we follow [25] and adopt a downsampling mean squared error (MSE) loss (denoted as ${{\cal L}_{DS}}$) for self-supervised training. This loss ensures that the enhanced super-resolution image remains globally consistent, on a pixel level, with the original low-resolution image. Let the downsampling function be ${H(\cdot)}$, and define ${J^{*}}$ as the median image, which must satisfy:
$H(Y)\approx J^{*}$ (6)
$J^{*}(u,v)=\operatorname{median}\{J^{(1)}(u,v),J^{(2)}(u,v),\ldots,J^{(M)}(u,v)\}$ (7)
where ${(u,v)}$ denotes the pixel coordinates in the image. Optimization is performed using the downsampling MSE loss:
$L_{down}=\|H(Y)-J^{*}\|_{2}^{2}$ (8)
However, because the downsampling loss primarily focuses on global pixel information, it tends to overly smooth high-frequency details, thereby weakening key features in target regions and affecting subsequent target detection. To address this issue, we introduce a local consistency loss and a gradient consistency loss to ensure that the enhanced super-resolution image preserves the local variation trends of the low-resolution image. By comparing the differences in adjacent pixel regions, we can effectively retain high-frequency details and avoid over-smoothing:
$L_{\mathrm{con}}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\left(\left|Y_{p}-Y_{q}\right|-\left|J_{p}^{*}-J_{q}^{*}\right|\right)$ (9)
$L_{\mathrm{grad}}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\left(\left|\nabla(Y_{p})-\nabla(Y_{q})\right|-\left|\nabla(J_{p}^{*})-\nabla(J_{q}^{*})\right|\right)$ (10)
[…]
The final loss function ${L_{\mathrm{total}}}$ is expressed as:
$L_{\rm total}=L_{\rm down}+\lambda_{\rm con}L_{\rm con}+\lambda_{\rm grad}L_{\rm grad}$ (11)
where ${\lambda_{\mathrm{con}}}$ and ${\lambda_{\mathrm{grad}}}$ are the hyperparameters that balance each component of the loss function.

### v3

# Reference-Free Enhancement of Forward-Looking Sonar Images: Bridging Cross-Modal Degradation Gaps Through Deformable Wavelet Scattering Transform and Multi-Frame Fusion

@@ Abstract
Enhancing forward-looking sonar images is crucial for effective underwater target detection. Existing deep learning approaches typically depend on supervised training using simulated data, but their performance and generalization are limited by the scarcity of real-world paired datasets. Although recent reference-free methods have mitigated reliance on paired data, they overlook the inherent cross-modal degradation gap between sonar and remote sensing images. Direct application of pretrained remote-sensing models to sonar data frequently yields overly smoothed images with lost details and inadequate brightness. To overcome this challenge, we propose a novel Deformable Wavelet Scattering Transform (WST) Feature Bridge, which adaptively projects sonar images into a multi-scale, translation-invariant feature domain, effectively narrowing the cross-modal degradation gap. Additionally, we develop a reference-free multi-frame fusion strategy that exploits complementary inter-frame information to suppress speckle noise and enhance target visibility without external ground-truth. Experiments conducted on three sonar datasets demonstrate that the proposed framework substantially surpasses existing reference-free methods in terms of noise suppression, edge preservation, and brightness enhancement, highlighting its practical effectiveness for underwater detection scenarios.
@@
Forward-Looking Sonar, reference-free, Multi-Frame Fusion, Cross-Modal Degradation Gap.
@@ I Introduction
Forward-looking sonar (FLS) provides unique underwater imaging capabilities, becoming essential in underwater target detection, environmental perception, and autonomous navigation tasks [1, 2, 3, 4, 5]. However, complex underwater environments significantly degrade sonar images due to multipath scattering, acoustic absorption by target materials, and reflections from environmental boundaries. Such effects typically introduce severe speckle noise, brightness inconsistency, and various imaging artifacts [6, 7, 8], hindering accurate target detection and reliable 3D reconstruction, thereby limiting practical deployment of FLS in real-world scenarios [9, 10, 11].
@@
Figure 1: (a) Bridging the Cross-Modal Gap. The degradation gap between multispectral remote sensing (RS) and forward-looking sonar (FLS) images is effectively narrowed by mapping both modalities into a unified feature space via our Deformable WST Feature Bridge; (b) Enhanced Results with Our Method. Compared to the overly smoothed baseline, our method preserves clear target contours and distinct foreground-background boundaries.
Current sonar image enhancement methods generally fall into two categories [12, 13, 14]: traditional filtering-based techniques and supervised deep-learning approaches. Traditional methods, including filtering, contrast adjustment, and grayscale transformations [15, 16, 17], often struggle to restore severely degraded sonar images effectively. Supervised deep-learning approaches [18, 19, 20, 21], while capable of achieving better enhancement, heavily depend on large-scale, high-quality paired datasets. Collecting such datasets in realistic underwater environments is notoriously challenging, which severely restricts their generalizability [22, 23]. Recent reference-free enhancement methods [24, 25] circumvent this dependence but have yet to address the intrinsic cross-modal degradation differences between forward-looking sonar and multispectral remote sensing images [26, 27, 28]. Although both modalities exhibit speckle noise, remote sensing images suffer from blur and low contrast due to uneven illumination and atmospheric attenuation under a pinhole camera model [29, 30], whereas sonar imaging is fundamentally distance-based, resulting in multipath scattering-induced brightness variations and weaker contrast [1, 2]. Consequently, directly applying pretrained remote-sensing models to sonar enhancement frequently produces overly smoothed outputs with loss of essential details (Fig. 1(a), (b)).
To address these limitations, we propose a novel reference-free sonar enhancement pipeline that integrates a Deformable Wavelet Scattering Transform (WST) Feature Bridge and a multi-frame fusion network into an end-to-end framework. Specifically, our Deformable WST Feature Bridge employs learnable perturbations on wavelet scales and orientations—carefully selected via ablation among multiple handcrafted features [31, 32, 33, 15, 34]—to adaptively map sonar images into a robust, translation-invariant feature domain, significantly narrowing the cross-modal degradation gap. Subsequently, our multi-frame fusion network leverages complementary information from consecutive sonar frames, naturally suppressing speckle noise and enhancing target-region brightness without relying on external high-quality ground truth. Unlike previous methods, our unified approach simultaneously optimizes adaptive feature extraction and multi-frame fusion, resulting in sharper contours, richer details, and enhanced brightness in sonar imagery.
The main contributions of this paper can be summarized as follows:
  • We propose a Deformable WST Feature Bridge that adaptively aligns sonar image features with pretrained remote-sensing distributions by embedding learnable offsets into wavelet scales and orientations, effectively reducing the cross-modal degradation gap.
  • We develop a reference-free multi-frame fusion framework for sonar image enhancement, utilizing inter-frame complementary information to suppress speckle noise and enhance target contrast without external reference data.
  • We construct a carefully annotated sonar image dataset featuring three material-diverse targets (rubber tire, metal torpedo, GRP frustum) captured along circular and linear trajectories, exhibiting distinct degradation patterns (multipath scattering, acoustic absorption). Precise pose annotations eliminate the need for image registration and enable rigorous benchmarking. Experiments demonstrate superior performance of our method over state-of-the-art baselines both visually and quantitatively.
Code and Dataset Release. The source code and the collected sonar dataset will be publicly released upon publication.
Figure 2: Overview of the proposed method. The model combines a Deformable WST Feature Bridge—with learnable wavelet scale and orientation offsets—and a multi-frame fusion network into one end-to-end pipeline, replacing the fixed WST extractor plus simple concat-and-$1\times 1$ convolution. This unified design closes the cross-modal degradation gap, reduces speckle noise, and boosts target brightness, producing images with crisper contours and richer edge detail.
@@ II-A Quality Degradation of Forward-Looking Sonar Images and Traditional Enhancement Methods
In addition, these methods incorporate various handcrafted feature extraction techniques, such as Histogram of Oriented Gradients (HOG), Canny edge detection (Canny), Haar-like features (HAAR), Gaussian Radial Edges (GRE) and Wavelet Scattering Transform (WST) [31, 32, 33, 15, 34]. By extracting key structural information, these techniques transform images from the pixel domain to a more robust feature space, improving image quality under noisy conditions. We argue that such feature-space transformations effectively reduce cross-modal degradation gaps and provide a practical framework for transferring enhancement models trained in other domains to sonar image processing tasks.
@@ II-B Deep-Learning-Based Sonar Image Enhancement Methods
To reduce reliance on paired data, methods combining traditional transform-domain techniques and deep learning have emerged. For instance, FDBANet [41] introduces a frequency-domain denoising attention module using Fast Fourier Transform (FFT) to convert sonar images to the frequency domain, filtering noise and enhancing high-frequency features. While this method works well for low-resolution images, its enhancement is limited due to the lack of a data-driven learning mechanism. reference-free methods [42, 43, 44, 45], however, does not require reference images and directly utilizes real low-resolution sonar images for training. The SEGSID framework [24] enhances local receptive fields and extracts global semantic information, overcoming traditional blind spot network (BSN) limitations in handling speckle noise with strong spatial correlation. Yet, it uses only single-image information, limiting its ability to enhance image resolution. In contrast, multi-frame fusion methods leverage complementary information across frames, naturally removing noise and improving resolution [25]. This method boosts the brightness of small target areas by fine-tuning a pre-trained remote sensing model but does not fully address the cross-modal degradation gap between multispectral remote sensing and sonar images [34]. In regions with complex structures or severe degradation, the enhanced image can become overly smooth, losing edge details and affecting subsequent target detection accuracy.
@@ III-A Framework Overview
As shown in Figure 2, our reference-free forward-looking sonar enhancement pipeline consists of two key modules: the Deformable WST Feature Bridge and the Multi-Frame Fusion Network. The Deformable WST Feature Bridge applies learnable offsets to wavelet scales and orientations, adaptively transforming each raw sonar frame into a robust, multi-scale feature representation that aligns with pretrained remote-sensing features. These per-frame feature tensors are then stacked into a sequence and passed to the Multi-Frame Fusion Network, which comprises an encoder, a fusion block, and a decoder. By integrating complementary information across consecutive frames, the fusion network suppresses speckle noise and reconstructs a high-resolution output. All components are learned end-to-end under a reference-free objective—combining downsampling consistency with local and gradient consistency losses—and leverage accurate pose metadata to eliminate any need for explicit registration.
III-B Deformable WST Feature Bridge
To effectively bridge the cross-modal degradation gap between multispectral remote sensing and forward-looking sonar imagery, we introduce the Deformable Wavelet Scattering Transform (WST) Feature Bridge. This module adaptively transforms raw sonar images into a robust, unified feature representation tailored specifically for sonar image enhancement.
Let $I\in\mathbb{R}^{H\times W}$ denote the input sonar image. We first obtain a set of handcrafted feature maps:
$G_{i}(I)=\Phi_{i}(I),\quad i\in\{\mathrm{HOG},\mathrm{Canny},\mathrm{GRE},\mathrm{Haar},\mathrm{WST}\},$ (1)
where each $\Phi_{i}(\cdot)$ applies the corresponding filter and $G_{i}(I)\in\mathbb{R}^{H\times W}$.
Based on our ablation study, the Wavelet Scattering Transform (WST) offers an optimal trade-off between denoising and detail preservation. To further adapt WST to sonar-specific patterns, we introduce learnable scale and orientation offsets $(\Delta j,\Delta\theta)$, enabling dynamic deformation of the original Morlet wavelets:
$\psi_{j,\theta}(x)\;\longmapsto\;\psi_{j+\Delta j,\,\theta+\Delta\theta}(x),\quad x\in\mathbb{R}^{2},$ (2)
where $j\in\{1,\dots,J\}$ and $\theta\in\{\theta_{1},\dots,\theta_{K}\}$, and each $\psi$ is centered on pixel coordinate $x$.
The first-order deformable scattering coefficients are defined by
$S_{1}(I;j,\theta)=\bigl{\lvert}\,I\ast\psi_{\,j+\Delta j,\,\theta+\Delta\theta}\bigr{\rvert}\ast\phi.$ (3)
[…]
Second-order deformable scattering coefficients further capture cross-scale interactions (computed only for $j_{1}<j_{2}$ or $\theta_{1}<\theta_{2}$ to avoid duplication):
$S_{2}(I;$ $\,j_{1},\theta_{1},j_{2},\theta_{2})$ (4)
$=\bigl{|}\,|I\ast\psi_{j_{1}+\Delta j_{1},\theta_{1}+\Delta\theta_{1}}|$
$\quad\;\ast\psi_{j_{2}+\Delta j_{2},\theta_{2}+\Delta\theta_{2}}\bigr{|}\ast\phi,$
Defining $S_{0}=I\ast\phi$, we concatenate all orders into the unified deformable WST feature tensor:
$G_{\mathrm{WST}}(I)=\mathrm{concat}\bigl{(}S_{0},\,S_{1},\,S_{2}\bigr{)},$ (5)
where ‘concat’ denotes channel-wise concatenation, yielding an output of size $H\times W\times\bigl{(}1+JK+\binom{JK}{2}\bigr{)}$. This tensor is then downsampled by a factor of 2 (bilinear interpolation) and channel-normalized via BatchNorm. Our implementation employs $J=3$ scales and $K=6$ orientations, generating $3\!\times\!6=18$ first-order and $\binom{18}{2}=153$ second-order channels. Crucially, all offsets $(\Delta j,\Delta\theta)$ are learned jointly with the multi-frame fusion network.
As illustrated in Fig. 1(a), the original forward-looking sonar image (red) and multispectral remote sensing image (green) occupy distinct pixel domains, creating a cross-modal degradation gap. Conventional approaches attempt to narrow this gap by concatenating handcrafted feature maps with the raw image and applying a fixed $1\times 1$ convolution. In contrast, our Deformable WST Feature Bridge (Fig. 1(b)) directly generates the unified, adaptively optimized feature tensor $G_{\mathrm{WST}}(I)$, shifting data into the yellow feature space. This dynamic feature adaptation aligns with the reference-free fusion objective, substantially reducing the degradation gap and improving the transferability of pretrained remote-sensing models for sonar image enhancement tasks.
@@ III-C Multi-Frame Image Fusion Network
Figure 3: Illustration of the multi-frame fusion network. The network recursively fuses aligned forward-looking sonar feature tensors in a pairwise manner until all $K$ frames are aggregated. A median reference image guides reference-free enhancement through downsampling and gradient consistency losses, enabling effective speckle noise suppression and high-resolution output (from $376\times 376$ to $800\times 800$).
Forward-looking sonar systems typically capture images at high frame rates, enabling complementary information across consecutive frames to naturally suppress speckle noise. In our pipeline, each raw frame $I^{(i)}\in\mathbb{R}^{H\times W}$ is first processed by the Deformable WST Feature Bridge to produce a unified feature tensor
$\tilde{J}^{(i)}=G_{\mathrm{WST}}\bigl{(}I^{(i)}\bigr{)},\quad i=1,\dots,K.$
We then form the input sequence
$J=\{\tilde{J}^{(1)},\tilde{J}^{(2)},\ldots,\tilde{J}^{(K)}\}.$
The network comprises three modules: encoding $E(\cdot)$, fusion $F(\cdot)$, and decoding $D(\cdot)$:
$Z$ $=F\bigl{(}E(J)\bigr{)},$ (6)
$Y$ $=D(Z),$ (7)
where $Y\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ is the super-resolved output.
For reference-free training, we adopt a downsampling mean squared error loss. Let $H(\cdot)$ denote bicubic downsampling by factor 2, and define the pixel-wise median reference
$J^{*}(u,v)=\operatorname{median}\bigl{\{}J^{(i)}(u,v)\ \big{|}\ i=1,\dots,K\bigr{\}}.$
The downsampling loss is
$L_{\rm down}\;=\;\bigl{\lVert}H(Y)-J^{*}\bigr{\rVert}_{2}^{2}.$ (8)
To preserve high-frequency details, we add a local consistency loss
$L_{\rm con}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\Bigl{(}\lvert Y_{p}-Y_{q}\rvert-\lvert J^{*}_{p}-J^{*}_{q}\rvert\Bigr{)}$ (9)
and a gradient consistency loss
$L_{\rm grad}=\frac{1}{P}\sum_{p=1}^{P}\sum_{q\in\sigma(p)}\Bigl{(}\lvert\nabla Y_{p}-\nabla Y_{q}\rvert-\lvert\nabla J^{*}_{p}-\nabla J^{*}_{q}\rvert\Bigr{)},$ (10)
where $p$ indexes local patches, $\sigma(p)$ its neighbors, and $\nabla$ the spatial gradient.
The overall objective is
$L_{\rm total}=L_{\rm down}\;+\;\lambda_{\rm con}L_{\rm con}\;+\;\lambda_{\rm grad}L_{\rm grad},$ (11)
with hyperparameters $\lambda_{\rm con},\lambda_{\rm grad}$ balancing each term.